### v1

# Displacement-Squeeze receiver for BPSK displaced squeezed vacuum states surpassing the coherent-states Helstrom bound under imperfect conditions

Abstract
We propose a displacement-squeeze receiver (DSR) for discriminating BPSK displaced squeezed vacuum states (S-BPSK). The receiver applies a displacement followed by a squeezing operation with the squeezing axis rotated by $\frac{\pi}{2}$, and performs photon-number-resolving detection with a MAP threshold decision. This processing effectively increases the distinguishability of the input states by elongating their distance in phase space and reducing their population overlap in Fock basis. We show that for all signal energy N, $P_{\text{err}}^{\text{DSR}}\in\left[P_{\text{HB}}^{\text{DSS}},2P_{\text{HB}}^{\text{DSS}}\right]$, under equal priors and ideal condition. In the low-energy regime, DSR beats the S-BPSK SQL at $N\approx 0.3$ and drops below the coherent-state BPSK (C-BPSK) Helstrom bound at $N\approx 0.4$, reaching $P_{\text{err}}^{\text{DSR}}<1\%$ near $N\approx 0.6$. Finally, we quantify performance under non-unit efficiency and dark counts, phase diffusion, and receiver thermal noise, with MAP threshold adaptation providing robustness across these nonidealities.
preprint: APS/123-QED
I Introduction
As an important topic in the field of quantum information, quantum-states discrimination directly affects the performance of quantum communication systems. To date, we have consistently regarded coherent states (CSs) as carriers for quantum communication and have proposed numerous excellent discrimination schemes, such as the Kennedy receiver[15], the optimized displacement receiver[24], multi-stage receiver[25, 14, 5, 18], the conditional pulse nulling receiver[32], and hybrid receiver [17, 1, 21, 33]. This is primarily because coherent states retain their coherence after passing through lossy bosonic channels, demonstrating that they can maintain state purity in most optical transmission media, including but not limited to free space and optical fibers.[8, 27, 16]. However, coherent states have inherent discrimination limitations: any two coherent states are non-orthogonal and exhibit overlap in phase space. Since the Helstrom Bound (HB) is closely related to the degree of overlap between quantum states[10, 11], coherent states have a theoretical lower bound on discrimination error probability.
In recent years, quantum-state discrimination schemes based on non-Poissonian photon statistics have attracted widespread attention[22, 13, 23, 6, 3, 26]. For example, Giovanni Chesi and Stefano Olivares proposed using squeezed vacuum states ($\hat{S}(r)\left|0\right\rangle$ where $\hat{S}(r),r\in\mathbf{R}_{+}$ is squeeze operator and $\left|0\right\rangle$ is vacuum state) as information carriers to construct a binary communication channel[6]. By applying a displacement operation $\hat{D}(\pm\alpha)$ to the squeezed vacuum state, two phase-opposite displaced squeezed vacuum states (DSSs) $\left|\pm\alpha,r\right\rangle$ can be generated. Then they encoded BPSK symbols into two phase-opposite DSSs (S-BPSK):
$\left\{\begin{aligned} \text{symbol}\ ^{\prime}0^{\prime}:\ &\left|\psi_{0}\right\rangle=\hat{D}(-\alpha)\hat{S}(r)\left|0\right\rangle=\left|-\alpha,r\right\rangle\\ \text{symbol}\ ^{\prime}1^{\prime}:\ &\left|\psi_{1}\right\rangle=\hat{D}(+\alpha)\hat{S}(r)\left|0\right\rangle=\left|+\alpha,r\right\rangle\end{aligned}\right$. (1)
This scheme significantly reduces the overlap between the carrier states by squeezing its position quadrature $\hat{X}$. Theoretical studies show that when the communication channel energy $N$ is fixed, not only the HB for S-BSPK can be significantly lower than that for BPSK coherent states (C-BPSK), but the error probability of homodyne detection (Standard Quantum Limit, SQL) for S-BSPK can even surpass the HB for C-BSPK at a low mean photon number of $N\approx 0.659$ [6].
Unlike the work of Giovanni Chesi and Stefano Olivares, this paper proposes a quantum-enhanced receiver for discriminating the S-BPSK signal, which is called as the Displacement-Squeeze Receiver (DSR). The DSR combines Gaussian operations (displacement and squeeze) and photon-number-resolving detection to approach the Helstrom bound for S-BPSK within a constant factor, while outperforming coherent-state receivers in the low-energy regime. Theoretical analysis shows that, for the same input mean photon number, the DSR can achieve an error probability lower than the Helstrom bound for C-BPSK channels in the low-photon-number regime. For example, when $N\approx 0.6$ the error probability falls below 1% in our simulations.
The structure of this paper is as follows: Section II reviews the fundamental theory of S-BPSK discrimination; Section III provides a detailed analysis of the principle and performance of the DSR; Section IV explores the performance of DSR under non-ideal conditions, including non-ideal PNR detectors (quantum efficiency $\eta$ and dark count rate $\nu$), phase diffusion noise in the channel, and receiver thermal noise; finally, the paper is summarized, and potential application scenarios are discussed.
II Channel model and performance benchmarks
In this section we summarize the signal model and the fundamental performance limits for binary phase-shift keyed (BPSK) communication based on displaced squeezed vacuum states, and we fix the notation used in the rest of the paper.
For binary discrimination between quantum states $\left|\psi_{0}\right\rangle$ and $\left|\psi_{1}\right\rangle$ with prior probabilities $P_{0}=P_{1}=\frac{1}{2}$, the minimum error probability (Helstrom bound) is given by [10, 6]:
$P_{\text{HB}}^{\text{DSS}}\left(\alpha,r\right)$ $=\frac{1}{2}\left(1-\sqrt{1-\left|\left\langle\psi_{0}|\psi_{1}\right\rangle\right|^{2}}\right)$ (2)
$=\frac{1}{2}\left(1-\sqrt{1-\exp\left(-4\alpha^{2}\text{e}^{2r}\right)}\right)$
As a practically relevant benchmark we also consider homodyne detection along the X quadrature, which attains the standard quantum limit (SQL) for BPSK under ideal conditions. Given the input is $\rho_{i}=\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|$, the conditional probability density of the homodyne outcome $x$ ,
$P\left(x|\rho_{i}\right)=\int_{-\infty}^{+\infty}\text{d}p\cdot W\left(x,p|\rho_{i}\right)$ (3)
$=\frac{\mathrm{e}^{r}}{\sqrt{\pi}}\,\exp\left\{-\,\mathrm{e}^{2r}\left[x-(-1)^{i+1}\sqrt{2}\alpha\right]^{2}\right\},\ i\in{0,1},$
which $W\left(x,p|\rho_{i}\right)$ is the Wigner function of $\rho_{i}$ (see Appendix). And the resulting error probability is
$P_{\text{SQL}}^{\text{DSS}}=\int_{0}^{+\infty}P\left(x|\rho_{0}\right)\mathrm{d}x=\frac{1}{2}\text{erfc}\left(\sqrt{2}\alpha\text{e}^{r}\right)$ (4)
To parameterize the contribution of squeezing, we introduce the squeezing fraction
$\beta:=\frac{\sinh^{2}(r)}{N}$ (5)
which quantifies the share of the total energy ($N=\left|\alpha\right|^{2}+\sinh^{2}(r)$) stored in squeezing [6]. Therefore, the parameter $\alpha$ and $r$ can be expressed as
$\left\{\begin{aligned} &\alpha=\sqrt{N\left(1-\beta\right)},\quad r=\sinh^{-1}\sqrt{N\beta}\\ &\alpha\text{e}^{r}=\sqrt{N\left(1-\beta\right)}\left(\sqrt{1+N\beta}+\sqrt{N\beta}\right)\end{aligned}\right$. (6)
Accordingly, both the Helstrom bound and the SQL for S-BPSK become functions of $\left(N,\beta\right)$, denoted $P_{\text{SQL}}^{\text{DSS}}\left(N,\beta\right)$, and $P_{\text{HB}}^{\text{DSS}}\left(N,\beta\right)$. The squeezing fraction can be optimized at fixed total energy. By minimizing $P_{\text{HB}}^{\text{DSS}}\left(N,\beta\right)$ over $\beta$ one obtains the optimal value [6]:
$\beta_{\text{opt}}(N)=\frac{N}{2N+1}$ (7)
Therefore, $\alpha^{2}\text{e}^{2r}=N(N+1)$. Both the Helstrom bound and the SQL for S-BPSK are rewritten as:
$P_{\text{SQL}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)=\frac{1}{2}\text{erfc}\left(\sqrt{2N\left(N+1\right)}\right)$ (8)
$P_{\text{HB}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)=\frac{1}{2}\left(1-\sqrt{1-\exp\left\{-4N(N+1)\right\}}\right)$ (9)
Figure 1: The ratios of $P_{\text{HB}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)$, $P_{\text{SQL}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)$ and $P_{\text{SQL}}^{\text{CS}}\left(N\right)$ to $P_{\text{HB}}^{\text{CS}}\left(N\right)$ vs. input state energy N. Here, $P_{\text{SQL}}^{\text{CS}}\left(N\right)=\frac{1}{2}\left[1-\mathrm{erf}\left(\sqrt{2\cdot N}\right)\right]$ and $P_{\text{HB}}^{\text{CS}}\left(N\right)=\frac{1}{2}\left[1-\sqrt{1-\exp\left(-4\cdot N\right)}\right]$.
When the DSSs are prepared with this choice, the S-BPSK Helstrom bound is significantly lower than that of C-BPSK at the same energy. In particular, for $N\approx 1$ the S-BPSK Helstrom bound already reaches the order of $10^{-4}$, whereas C-BPSK requires approximately twice the energy to achieve a comparable error probability. Moreover, in the low-photon number regime the S-BPSK SQL can fall below the Helstrom bound of C-BPSK, with the crossover occurring around $N\approx 0.659$ (see Fig. 1). At $N=1.0$, the HB for S-BPSK surpasses the SQL for C-BPSK by a remarkable 17.39dB, and the SQL for S-BPSK also is 2.94dB lower. In the subsequent analyses, all DSSs employ the optimal squeezing fraction $\beta_{\text{opt}}$.
III Quantum-enhanced Receiver for discriminating BPSK Displaced Squeezed Vacuum States
We propose a quantum-enhanced receiver for S-BPSK discrimination, referred to as the displacement–squeeze receiver (DSR). Its schematic is shown in Fig. 2. Compared with conventional displacement receivers, the DSR introduces an additional squeezing operation. The squeezing strength matches that used at the transmitter, whereas the squeezing axis is rotated by $\pi/2$ in phase space.
Figure 2: Schematic of the proposed DSR. The input state $\rho_{i}$ is transformed to $\zeta_{i}$ via a displacement followed by a squeezing operation, $U=S(-r)D(\alpha)$, and is then measured by a photon-number-resolving (PNR) detector with outcome $n$. The final decision is made according to the maximum a posteriori (MAP) criterion.
In the DSR, the input state $\rho_{i}=\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|$ first interferes with a local oscillator $\left|\alpha\right\rangle$ on a highly transmissive beam splitter ($\tau\to 1$), implementing the displacement operation $D(\alpha)$. This maps the BPSK pair $\left\{\left|-\alpha,r\right\rangle,\left|+\alpha,r\right\rangle\right\}$ to an OOK-like pair $\left\{\left|0,r\right\rangle,\left|+2\alpha,r\right\rangle\right\}$. Subsequently, a squeezing operation S(r)S(-r) is applied, squeezing the $P$ quadrature. The processed state is denoted by
$\zeta_{i}=U\rho_{i}U^{\dagger},\quad U=S(-r)D(\alpha).$ (10)
A PNR detector then measures $\zeta_{i}$ and outputs the photon number $n$, based on which the receiver makes a MAP decision.
Figure 3: Phase-space representations (a–c) and Fock-basis populations (d–f) of the input states $\rho_{i}$ in the DSR for mean photon number $N=1.0$ and squeezing fraction $\beta=\beta_{\text{opt}}$. Panels (a) and (d) show the input states $\rho_{i}$; panels (b) and (e) show the states after the displacement operation $D\rho_{i}D^{\dagger}$; and panels (c) and (f) show the states after the full DSR transformation $\zeta_{i}=U\rho_{i}U^{\dagger}$.
We next provide an intuitive explanation of the DSR advantage. Figs. 3(a)–(c) illustrate how the phase-space distributions evolve under the displacement and squeezing operations for $N=1.0$ and $\beta=\beta_{\text{opt}}$. The two input states $\rho_{0}$ and $\rho_{1}$ are separated by $\Delta=2\alpha$ in phase space. After the displacement operation, the two states become a squeezed vacuum state and a displaced squeezed vacuum state, respectively (Fig. 3(b)):
$\left\{\begin{aligned} &D(\alpha)\cdot\left|\psi_{0}\right\rangle=S(r)\left|0\right\rangle\\ &D(\alpha)\cdot\left|\psi_{1}\right\rangle=D(2\alpha)S(r)\left|0\right\rangle\end{aligned}\right$. (11)
The displacement does not change the separation $\Delta=2\alpha$ between the state centroids in phase space; however, it leads to distinct Fock-basis populations for $\rho_{0}$ and $\rho_{1}$ (see Figs. 3(d) and (e)).
After the subsequent squeezing operation, the two states evolve into the vacuum state and a coherent state, respectively (Fig. 3(c)):
$\left\{\begin{aligned} &U\cdot\left|\psi_{0}\right\rangle=\left|0\right\rangle\\ &U\cdot\left|\psi_{1}\right\rangle=S(-r)D(2\alpha)S(r)\left|0\right\rangle=D(\gamma)\left|0\right\rangle=\left|\gamma\right\rangle\end{aligned}\right$. (12)
Here,
$\gamma$ $=2\alpha\cdot\cosh(r)+2\alpha^{*}\cdot\sinh(r)$ (13)
$=2\alpha\cdot\mathrm{e}^{r}>2\alpha,$
where the last equality assumes $\alpha$ is real (i.e., the phase reference is chosen such that $\alpha\in\mathbb{R}$). Therefore, the S-BPSK input $\rho_{i}\in\left\{\left|-\alpha,r\right\rangle,\left|+\alpha,r\right\rangle\right\}$ is converted into a coherent-state OOK pair $\zeta_{i}\in\left\{\left|0\right\rangle,\left|\gamma\right\rangle\right\}$. Importantly, the phase-space separation increases from $2\alpha$ to $\gamma=2\alpha\,\mathrm{e}^{r}$, enhancing distinguishability for subsequent PNR detection. By comparison, a coherent-state displacement receiver (without squeezing) yields a separation $\Delta_{\mathrm{C}}=2\sqrt{N}$, whereas the DSR produces $\gamma=2\sqrt{N(N+1)}$.
Figs. 3(d)–(f) show the corresponding changes in Fock-basis populations. After the squeezing operation (Fig. 3(f)), the population associated with $\rho_{1}$ shifts toward higher photon numbers, while the population associated with $\rho_{0}$ concentrates near $n=0$. Consequently, the overlap between the two photon-number distributions is reduced, facilitating discrimination with a PNR detector.
III.1 PNR detection model and MAP decision rule
We now incorporate finite PNR resolution. A PNR detector with resolution $M$ is denoted as PNR(M)(M) and is described by the POVM elements
$\Pi_{n}=\left|n\right\rangle\left\langle n\right|,\quad n=0,1,.,M-1$ (14)
$\Pi_{M}=\mathbf{I}-\sum_{i=0}^{M-1}\Pi_{i}$
where $\mathbf{I}$ is the identity operator in the Fock basis. For a measured coherent state $\rho$ with mean photon number $\mu$, the conditional probability of obtaining outcome $n$ is
$q_{n}^{M}\left(\mu\right)=\mathbf{Tr}\left\{\Pi_{n}\cdot\rho\right\}=\left\{\begin{aligned} &\mathrm{e}^{-\mu}\cdot\frac{\mu^{n}}{n!},\quad n\leq M-1\\ &1-\mathrm{e}^{-\mu}\cdot\sum_{j=0}^{M-1}\frac{\mu^{j}}{j!},\quad n=M\end{aligned}\right$. (15)
Accordingly, the conditional probability of outcome $n$ given that symbol $i$ was sent is
$P(n|i)=\mathbf{Tr}\left(\Pi_{n}\cdot\zeta_{i}\right)=\left\{\begin{aligned} &q_{n}^{M}(0),\ i=0\\ &q_{n}^{M}\left(|\gamma|^{2}\right),\ i=1\end{aligned}\right$. (16)
The MAP decision rule for binary discrimination can be equivalently expressed as an optimal threshold test on $n$ [28, 20], summarized in Table 1. The corresponding threshold $n_{\text{th}}^{*}$ satisfies Eq. (17).
$P_{0}\cdot P(n|0)\gtreqless P_{1}\cdot P(n|1),\ \text{where}\ n\lesseqgtr n_{\text{th}}^{*}$ (17)
$\Rightarrow n_{th}^{*}=\min\left\{\lceil n\rceil,\ \text{M}\right\}$
where, $\lceil*\rceil$ denotes the ceiling function (rounding up); Under ideal and equal-prior probability conditions, one obtains $n_{\text{th}}^{*}=1$.
Table 1: Threshold decision strategy of the DSR. If the PNR(M)(M) outcome satisfies $n<n_{\text{th}}^{*}$, decide symbol ‘0’; otherwise decide symbol ‘1’.
The outcomes of PNR(M) Decision
$n\geq n_{\text{th}}^{*}$ symbol ’1’
$n<n_{\text{th}}^{*}$ symbol ’0’
III.2 Error probability under ideal conditions
Now we analyze the error probability of the DSR. Assuming the optimal threshold value is $n_{th}^{*}$, the threshold decision strategy (see Table. 1) can be expressed using POVM operators:
$\text{symbol}\ ^{\prime}0^{\prime}:\quad\Pi_{\text{off}}=\sum_{n=0}^{n_{th}^{*}-1}\Pi_{n}=\sum_{n=0}^{n_{th}^{*}-1}\left|n\right\rangle\left\langle n\right|$ (18)
$\text{symbol}\ ^{\prime}1^{\prime}:\quad\Pi_{\text{on}}=\mathbf{I}-\Pi_{\text{off}}$
The average error probability is then
$P_{\text{err}}^{\text{DSR}}$ $=1-\frac{1}{2}\left[P(\text{C}|0)+P(\text{C}|1)\right]$ (19)
$=1-\frac{1}{2}\left[\mathbf{Tr}\left(\Pi_{\text{off}}\cdot\zeta_{0}\right)+\mathbf{Tr}\left(\Pi_{\text{on}}\cdot\zeta_{1}\right)\right]$
where $P(\text{C}|i)$ denotes the probability of a correct decision conditioned on sending symbol $i$. Under ideal conditions, $n_{\text{th}}^{*}=1$, hence $\Pi_{\text{off}}=\left|0\right\rangle\left\langle 0\right|,\Pi_{\text{on}}=\mathbf{I}-\left|0\right\rangle\left\langle 0\right|$. The DSR error probability becomes
$P_{\text{err}}^{\text{DSR}}$ $=1-\frac{1}{2}\left[q_{0}^{M}(0)+1-q_{0}^{M}\left(|\gamma|^{2}\right)\right]$ (20)
$=\frac{1}{2}\exp\left\{-4|\alpha|^{2}\mathrm{e}^{2r}\right\}=\frac{1}{2}\exp\left\{-4N\left(N+1\right)\right\}$
Eq. (20) shows that, under ideal conditions, $P_{\text{err}}^{\text{DSR}}$ is independent of the PNR resolution $M$. Therefore, in the numerical results of Fig. 4, we consider PNR(1), i.e., a single-photon detector (SPD).
For S-BPSK, the Helstrom bound in Eq. (9) and the DSR error probability in Eq. (20) satisfy
$\frac{P_{\text{err}}^{\text{DSR}}}{P_{\text{HB}}^{\text{DSS}}}=\frac{\exp\left\{-4N\left(N+1\right)\right\}}{1-\sqrt{1-\exp\left\{-4N(N+1)\right\}}}$ (21)
In the low-energy limit $N\to 0$, both $P_{\text{err}}^{\text{DSR}}$ and $P_{\text{HB}}^{\text{DSS}}$ approach $0.5$ and the ratio tends to $1$, corresponding to random guessing. In the high-energy regime $N\gg 1$, the ratio approaches $2$ (i.e., $3$ dB), yielding
$P_{\text{err}}^{\text{DSR}}\simeq 2P_{\text{HB}}^{\text{DSS}}$ (22)
Although the DSR remains suboptimal with respect to the Helstrom bound of S-BPSK and satisfies $P_{\text{err}}^{\text{DSR}}\in\left[P_{\text{HB}}^{\text{DSS}},2P_{\text{HB}}^{\text{DSS}}\right]$ for all $N$, it still provides a substantial advantage over C-BPSK channels. In particular, for the same mean photon number $N$, the proposed DSR can achieve an error probability below the Helstrom bound $P_{\text{HB}}^{\text{CS}}$ of a C-BPSK channel; for example, around $N\approx 0.4$ in Fig. 4 the DSR performance lies between the DSS Helstrom bound and the coherent state Helstrom bound.
Figure 4: Performance of the DSR with an SPD under ideal conditions. (a) Error probability of the DSR. (b) Ratio (in dB) of the benchmarks $P_{\text{HB}}^{\text{CS}}$, $P_{\text{SQL}}^{\text{CS}}$, $P_{\text{HB}}^{\text{DSS}}$, and $P_{\text{SQL}}^{\text{DSS}}$ relative to $P_{\text{err}}^{\text{DSR}}$.
Figure 4(a) shows the DSR error probability as a function of the mean photon number. The DSR surpasses the C-BPSK SQL, the S-BPSK SQL, and the C-BPSK Helstrom bound at $N\approx 0.21$, $0.30$, and $0.40$, respectively. As $N$ increases, the performance advantage over these benchmarks grows. At $N=1.0$, the DSR achieves gains of $21.3$ dB, $11.4$ dB, and $14.4$ dB relative to the C-BPSK SQL, the S-BPSK SQL, and the C-BPSK Helstrom bound, respectively, while remaining $3$ dB above the S-BPSK Helstrom bound, as summarized in Fig. 4(b). These results demonstrate that the DSR provides strong performance over a broad photon-number regime.
The above analysis assumes ideal, noiseless conditions. In the next section, we examine the impact of non-ideal factors on receiver performance, including imperfect photon detection, phase diffusion, and thermal noise.
IV The performance of DSR under realistic conditions
IV.1 The quantum efficiency $\eta$ and the dark count rate $\nu$ of PNR(M)
The realistic factors of PNR detectors include quantum efficiency $\eta$ and dark count rate $\nu$. Quantum efficiency essentially corresponds to photon loss during detection, while dark count rate fundamentally arises from random triggering events caused by environmental noise [12, 7]. Under ideal conditions, quantum efficiency $\eta=1$ and dark count rate $\nu=0$. If we simultaneously consider imperfect quantum efficiency $\eta<1$ and dark count rate $\nu>0$, the POVM operators for PNR(M) (Eq. (14)) are rewritten as [2]:
$\Pi_{n}=\mathrm{e}^{-\nu}\sum_{l=0}^{n}\sum_{k=n-l}^{\infty}\frac{\nu^{l}}{l!}\mathrm{C}_{n-l}^{k}\cdot\eta^{n-l}(1-\eta)^{k-(n-l)}\left|k\right\rangle\left\langle k\right|,\ n=0,...,M$ (23)
At this point, the probability of PNR(M) detecting $n$ photons (Eq. (15)) is rewritten as:
$q_{n}^{M}\left(\mu,\eta,\nu\right)=\left\{\begin{aligned} &\mathrm{e}^{-(\eta\mu+\nu)}\cdot\dfrac{\left(\eta\mu+\nu\right)^{n}}{n!},\quad n\leq M-1\\ &1-\mathrm{e}^{-(\eta\mu+\nu)}\cdot\sum_{j=0}^{M-1}\dfrac{\left(\eta\mu+\nu\right)^{j}}{j!},\quad n=M\end{aligned}\right$. (24)
By substituting into Eq. (17), the optimal threshold $n_{th}^{*}$ is obtained:
$n_{\text{th}}^{*}=\min\left\{\left\lceil\frac{\eta|\gamma|^{2}}{\ln\left(\eta|\gamma|^{2}+\nu\right)-\ln\nu}\right\rceil,M\right\}=\min\left\{\left\lceil\frac{\eta\cdot 4N(N+1)}{\ln\left(\eta\cdot 4N(N+1)+\nu\right)-\ln\nu}\right\rceil,M\right\}$ (25)
If only consider quantum efficiency ($\eta<1,\ \nu=0$), the optimal threshold $n_{th}^{*}$ remains 1 ($\frac{\eta|\gamma|^{2}}{\ln\left(\eta|\gamma|^{2}+\nu\right)-\ln\nu}\to 0^{+},\ \left\lceil\frac{\eta|\gamma|^{2}}{\ln\left(\eta|\gamma|^{2}+\nu\right)-\ln\nu}\right\rceil\to 1$). Then, the error probability expression (Eq. (20)) is rewritten as:
$P_{\text{err}}^{\text{DSR}}(N,\eta)$ $=1-\frac{1}{2}\left[q_{0}^{M}(0,\eta)+1-q_{0}^{M}\left(|\gamma|^{2},\eta\right)\right]$ (26)
$=\frac{1}{2}\exp\left\{-4|\alpha|^{2}\cdot\mathrm{e}^{2r}\cdot\eta\right\}=\frac{1}{2}\exp\left\{-4N\left(N+1\right)\cdot\eta\right\}$
We can see that under non-ideal quantum efficiency condition, the resolution M of the PNR(M) in the DSR does not affect the error probability. Fig. 5(a) illustrates the variation of $P_{\text{err}}^{\text{DSR}}(N,\eta)$ (Eq. (26)) with respect to the average signal photon number N for $\eta$ = 1.0, 0.9, and 0.8. As shown, the three curves surpass the SQL of S-BPSK at N = 0.30, 0.39, and 0.56, respectively. Moreover, as N increases, the performance advantage of $P_{\text{err}}^{\text{DSR}}(N,\eta)$ over the SQL becomes more pronounced. At N = 1.0, the ratios of the three curves to the SQL are -4.5 dB, -7.97 dB, -11.44 dB, respectively.
Likewise, if only consider dark counts ($\eta=1,\ \nu>0$), the optimal threshold $n_{th}^{*}$ reads:
$n_{\text{th}}^{*}=\min\left\{\left\lceil\frac{|\gamma|^{2}}{\ln\left(|\gamma|^{2}+\nu\right)-\ln\nu}\right\rceil,M\right\}=\min\left\{\left\lceil\frac{4N(N+1)}{\ln\left(4N(N+1)+\nu\right)-\ln\nu}\right\rceil,M\right\}$ (27)
As shown, under non-ideal dark count conditions, $n_{\text{th}}^{*}$ is not always unity and varies with the average photon number (see Fig. 5(c)). In other words, in non-ideal dark count condition, the error probability of the DSR is correlated with the resolution of the PNR.
Figure 5: Performance of the DSR with an non-ideal PNR(M) which quantum efficiency $\eta<1$ and dark count $\nu>0$. (a) and (b) Error probability of the DSR. (c) Optimal threshold $n_{\text{th}}^{*}$. (d) Ratio (in dB) of the benchmark $P_{\text{SQL}}^{\text{DSS}}$ relative to $P_{\text{err}}^{\text{DSR}}$.
Fig. 5(b) shows the error probability as a function of the average signal photon number under two scenarios: considering only non-ideal dark counts, and considering both non-ideal detection efficiency and dark counts. In the presence of dark counts, the error probability exhibits an oscillatory pattern, which results from the continuous increase of the optimal threshold. This optimal threshold is limited by the resolution of the PNR detector; once it reaches the maximum resolvable photon number, the error probability no longer decreases with increasing average photon number but instead saturates. Furthermore, we observe that for curves with the same dark count rate but different detection efficiencies, the saturated error probabilities coincide. In contrast, for curves with the same detection efficiency but different dark count, a lower dark count leads to a lower saturated error probability.
We therefore conclude that the saturated error probability is determined primarily by the dark count, and its effect cannot be mitigated by increasing the signal strength. On the other hand, the performance degradation caused by non-ideal detection efficiency can be compensated by using a stronger signal. For practical PNR detectors, high detection efficiency and low dark count rate generally exhibit a trade-off relationship; improving the detection efficiency typically increases the dark count [30, 19]. Hence, in actual experiments, it can be beneficial to moderately reduce the detection efficiency in order to obtain a lower dark count, while compensating for the resultant loss in sensitivity by increasing the signal power.
In addition, a high dark count forces the optimal threshold to increase frequently until it reaches the resolution limit of the PNR detector, thereby driving the DSR scheme into saturation (see Fig. 5(c)). Consequently, a configuration with lower detection efficiency and lower dark count can alleviate the requirement for a PNR detector with very high resolution. Fig. 5(d) plots the ratio (in dB) of the error probability to the SQL of S-BPSK. It can be seen that even with both non-ideal detection efficiency and dark counts, the DSR scheme still surpasses the SQL of S-BPSK. For example, when $\eta=0.9$ and $\nu=10^{-2}$, the error probability of DSR with PNR(M) is lower than the SQL by -0.96dB (M=1), -3.07dB (M=2), -6.04dB (M=3), -9.58dB (M=4), and -13.52dB (M=5).
IV.2 Phase Diffuse Noise in communication channel
Phase diffusion noise is a common communication channel noise that increases the overlap between states (S-BPSK), making them more difficult to distinguish [20, 6]. A schematic of a phase diffusion noise channel is shown in Fig. 6.
Figure 6: Schematic diagram of phase diffusion noise channel. The signal state $\rho_{i}$ (pure state) is transformed into the state $\rho_{i}^{pd}$ (mixed state) in the phase diffuse communication channel.
After DSSs are contaminated by phase diffusion noise in the channel, they evolve from pure states to mixed states and can only be expressed using density operators. Let the uncontaminated DSS be denoted as $\rho_{i}=\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|$. The expression for the contaminated quantum state $\rho_{i}^{\text{pd}}$ is then:
$\rho_{i}^{\text{pd}}$ $=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g_{\sigma}(\phi)\cdot\rho_{i,\phi}\right\}$ (28)
$=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g_{\sigma}(\phi)\cdot R_{\phi}\rho_{i}R_{\phi}^{\dagger}\right\}$
Here, $g_{\sigma}(\phi)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left\{-\frac{\phi^{2}}{2\sigma^{2}}\right\}$ is a Gaussian function with mean 0 and standard deviation $\sigma$, which quantifies the noise strength. The rotation operator $R_{\phi}=\mathrm{e}^{-j\phi a^{\dagger}a}$ rigidly rotates the entire Wigner function distribution of the quantum state by an angle $\phi$ in phase space. We denote the rotated quantum state as
$\rho_{i,\phi}=\left|\psi_{i,\phi}\right\rangle\left\langle\psi_{i,\phi}\right|$ (29)
$\left|\psi_{i,\phi}\right\rangle=R_{\phi}\left|\psi_{i}\right\rangle=D\left((-1)^{i+1}\alpha\mathrm{e}^{-j\phi}\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$
Thus, phase diffusion noise applies a Gaussian-distributed random phase rotation to the input signal, resulting in a completely positive map that transforms the input Gaussian state $\rho_{i}$ into a non-Gaussian state $\rho_{i}^{\text{pd}}$. The phase-space evolution of quantum state $\rho_{i}$ under phase diffusion noise is illustrated in Fig. 7, with signal mean photon number $N=1.0$ and a noise strength of $\sigma=0.5$.
Figure 7: Phase-space representations of S-BPSK signal (mean photon number $N=1.0$ and phase diffusion strength $\sigma=0.5$). (a) DSSs $\rho_{i}$ without phase diffusion noise; (b) states $\rho_{i}^{\text{pd}}$ with phase diffusion noise; (c) states $\zeta_{i}$ after DSR processing.
For the DSR receiver, since the input state $\rho_{i}^{\text{pd}}$ is no longer Gaussian, the quantum state $\zeta_{i}$ after displacement-squeezing processing and prior to PNR detection also ceases to be Gaussian.
$\zeta_{i}$ $=U\rho_{i}^{\text{pd}}U^{\dagger}$ (30)
$=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g(\phi)\cdot UR_{\phi}\rho_{i}R_{\phi}^{\dagger}U^{\dagger}\right\}$
$=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g(\phi)\cdot\left|\psi_{i,\phi}^{\prime}\right\rangle\left\langle\psi_{i,\phi}^{\prime}\right|\right\}$
where,
$\left|\psi_{0,\phi}^{\prime}\right\rangle=U\cdot\left|\psi_{0,\phi}\right\rangle=S\left(-r\right)D\left(-\alpha\mathrm{e}^{-j\phi}+\alpha\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$ (31)
$\left|\psi_{1,\phi}^{\prime}\right\rangle=U\cdot\left|\psi_{1,\phi}\right\rangle=S\left(-r\right)D\left(+\alpha\mathrm{e}^{-j\phi}+\alpha\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$
Therefore, when measuring the quantum state $\zeta_{i}$ with PNR(M), the probability of obtaining a detection result of $n$ photons is:
$q_{n}^{M}$ $\left(\zeta_{i}\right)=\mathbf{Tr}\left\{\Pi_{n}\cdot\zeta_{i}\right\}$ (32)
$=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot\mathbf{Tr}\left\{\Pi_{n}\cdot\left|\psi_{i,\phi}^{\prime}\right\rangle\left\langle\psi_{i,\phi}^{\prime}\right|\right\}$
By substituting into Eq. (17) and Eq. (19), we obtain the optimal threshold and the error probability of the DSR under phase diffusion noise, respectively.
Figure 8: Simulation results of DSR under phase diffusion noise ($\sigma=0.1$): (a) Error probability, (b) Ratio (in dB) with a benchmark of $P_{\text{SQL}}^{\text{DSS}}$ (Eq. (45)), (c) Optimal threshold $n_{th}^{*}$ and (d) Fock-basis population of measured states $\zeta_{i}=U\rho_{i}^{\text{pd}}U^{\dagger}$ by PNR(M), which $N=1.0$. Note: the SQL and HB of S-BPSK under phase diffusion condition is rewritten as Eq. (45) and Eq. (47).
Fig. 8(c) shows the variation of the optimal threshold $n_{\text{th}}^{*}$ for DSR with the mean photon number under a phase diffusion of $\sigma=0.1$. In the absence of resolution limits from the PNR detector, $n_{\text{th}}^{*}$ increases in steps of 2, which differs from the behaviors previously analyzed for dark counts and thermal noise. This is attributed to the fact that, in the presence of phase diffusion noise, DSR cannot transform $\rho_{0}^{\text{pd}}$ into the vacuum state or a coherent state with a very small mean photon number (see Fig. 7(c)). After DSR processing, the Fock-basis population of the resulting state $\zeta_{i}$ shows a feature reminiscent of squeezed states: the populations of certain even photon numbers become higher than those of adjacent odd photon numbers, as illustrated in Fig. 8(d).
Fig. 8(c) and (b) depict the performance of DSR as a function of the mean signal photon number under a phase diffusion strength of $\sigma=0.1$. The results show that DSR can surpass the SQL of the S-BPSK scheme within a certain photon-number range ($N\in\left(0.28,0.93\right)$), achieving a maximum advantage of 3.32 dB beyond the SQL. However, at high photon numbers, DSR no longer exceeds the SQL. In particular, when the optimal threshold reaches the maximum resolution limit of the PNR detector, the error probability does not saturate but instead increases with the mean photon number. Furthermore, we observe that a PNR detector with a resolution of 2n (even) provides only marginal improvement in DSR performance compared to one with a resolution of 2n-1; no pronounced oscillatory drop in error probability is evident. This behavior is consistent with the variation in the optimal threshold, as shown in panel (c).
IV.3 Thermal noise in the receiver
After contamination by thermal noise, the optical states are generally mixed and should be described by density operators. Following the Glauber–Sudarshan representation, the impact of thermal background at the receiver can be modeled as a classical, Gaussian-distributed random displacement in phase space, whose statistics are independent of the input state [28, 31, 29, 4]. Accordingly, for an arbitrary input state $\rho_{i}$, the thermal-noise channel can be written as
$\rho_{i}^{\mathrm{th}}=\Phi_{n_{t}}(\rho_{i})=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\,f_{n_{t}}(\lambda)\,D(\lambda)\rho_{i}D^{\dagger}(\lambda),$ (33)
where $D(\lambda)=\exp(\lambda a^{\dagger}-\lambda^{\ast}a)$ denotes the displacement operator and
$f_{n_{t}}(\lambda)=\frac{1}{\pi n_{t}}\exp\!\left(-\frac{|\lambda|^{2}}{n_{t}}\right)$ (34)
is the circularly symmetric complex Gaussian distribution. Here $n_{t}$ is the mean photon number of the thermal background mode, which under thermal equilibrium is
$n_{t}=\left[\exp\!\left(\frac{h\nu}{kT_{0}}\right)-1\right]^{-1},$ (35)
with $h$ the Planck constant, $k$ the Boltzmann constant, $\nu$ the optical frequency, and $T_{0}$ the receiver temperature. In practical fiber-optic communication systems, the receiver-originated background often dominates over the channel-induced thermal noise [9].
Consistent with [31, 29], we assume that the thermal-noise contamination occurs immediately before the PNR detector, as illustrated in Fig. 9.
Figure 9: Schematic of the DSR under receiver thermal noise. The processed states $\zeta_{i}$ impinging on the PNR(M) are contaminated by receiver thermal noise and become $\zeta_{i}^{\prime}$.
Remark (coherent state $\Rightarrow$ displaced thermal state): Under the channel $\Phi_{n_{t}}$ in Eq. (33), a coherent-state input becomes a displaced thermal state. Specifically, using the Weyl relation $D(\lambda)D(\alpha)=e^{(\lambda\alpha^{\ast}-\lambda^{\ast}\alpha)/2}D(\lambda+\alpha)$, one verifies the displacement covariance
$\Phi_{n_{t}}\!\big(D(\alpha)\rho D^{\dagger}(\alpha)\big)=D(\alpha)\Phi_{n_{t}}(\rho)D^{\dagger}(\alpha).$ (36)
Moreover, by definition of Eq. (33), $\Phi_{n_{t}}(|0\rangle\langle 0|)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\,f_{n_{t}}(\lambda)\,|\lambda\rangle\langle\lambda|$ is exactly a thermal state with mean photon number $n_{t}$, denoted as $\rho_{\mathrm{th}}(n_{t})$. Therefore, for $|\alpha\rangle=D(\alpha)|0\rangle$,
$\Phi_{n_{t}}\!\big(|\alpha\rangle\langle\alpha|\big)=\Phi_{n_{t}}\!\big(D(\alpha)|0\rangle\langle 0|D^{\dagger}(\alpha)\big)=D(\alpha)\rho_{\mathrm{th}}(n_{t})D^{\dagger}(\alpha).$ (37)
Consequently, the states $\zeta_{i}$ measured by PNR(M), originally $\left|0\right\rangle$ and $\left|2\alpha\mathrm{e}^{r}\right\rangle$, are transformed into
$\zeta_{0}^{\prime}$ $=\Phi_{n_{t}}\!\big(|0\rangle\langle 0|\big)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\,f_{n_{t}}(\lambda)\,\left|\lambda\right\rangle\left\langle\lambda\right|,$ (38)
$\zeta_{1}^{\prime}$ $=\Phi_{n_{t}}\!\big(|2\alpha\mathrm{e}^{r}\rangle\langle 2\alpha\mathrm{e}^{r}|\big)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\,f_{n_{t}}(\lambda)\,\left|\lambda+2\alpha\mathrm{e}^{r}\right\rangle\left\langle\lambda+2\alpha\mathrm{e}^{r}\right|.$
Equivalently, by Eq. (37), one may write $\zeta_{0}^{\prime}=\rho_{\mathrm{th}}(n_{t})$ and $\zeta_{1}^{\prime}=D(2\alpha\text{e}^{r})\rho_{\mathrm{th}}(n_{t})D^{\dagger}(2\alpha\text{e}^{r})$.
Thus, under receiver thermal noise, $\zeta_{0}^{\prime}$ and $\zeta_{1}^{\prime}$ are classical mixtures of coherent states. The probability of detecting $n$ photons with PNR(M) is
$q_{n}^{M}(\zeta_{i}^{\prime})=\mathrm{Tr}\!\left(\Pi_{n}\,\zeta_{i}^{\prime}\right)=\int_{\mathbb{C}}\mathrm{d}^{2}\lambda\,f_{n_{t}}(\lambda)\,q_{n}^{M}[\mu_{i}(\lambda)],$ (39)
where $\Pi_{n}$ is the POVM element corresponding to outcome $n$, and $\mu_{0}(\lambda)=|\lambda|^{2}$, $\mu_{1}(\lambda)=|\lambda+2\alpha\mathrm{e}^{r}|^{2}$. Substituting Eq. (39) into Eq. (17) and Eq. (19) yields the optimal threshold $n_{\text{th}}^{*}$ and the error probability $P_{\text{err}}^{\text{DSR}}$ of the DSR in the presence of receiver thermal noise, respectively.
Figure 10: Simulation results of DSR with PNR(M) under thermal noise condition: (a) Error probability ($n_{t}=0.001$), (b) Error probability ($M=10$), (c) Ratio (in dB) of the benchmark $P_{\text{err}}^{\text{DSS}}$ ($M=10$), (d) Optimal threshold ($M=10$), (e) Fock-basis population of $\zeta_{i}^{\prime}$ (the input energy $N=1.0$ and $n_{t}=0.001$), and (f) Maximum tolerable thermal noise $n_{t}^{\text{max}}$ for different photon number resolution M.
Fig. 10(a) and (b) depict the error probability of the DSR as a function of the mean signal photon number, under different PNR detector resolutions and various thermal noise strengths. The error probability curves exhibit an oscillatory pattern and decrease overall with increasing signal energy, which is a consequence of the continuous adjustment of the optimal threshold $n_{\text{th}}^{*}$ (see panel (d)). This trend continues until $n_{\text{th}}^{*}$ reaches the maximum resolution of the PNR detector, after which the error probability saturates and no longer varies with the mean photon number. Fig. 10(c) shows the ratio (in dB) of the DSR error probability to the SQL of standard BPSK versus the mean signal photon number under different noise strengths. It can be observed that the DSR can still surpass the SQL at certain noise levels. For instance, when $n_{t}=10^{-3}$, the error probability of the DSR with a PNR(M) detector is lower than the SQL by -6.34dB (M=1), -13.43dB (M=2), -20.79dB (M=3), and -28.27dB (M=4).
Furthermore, we note that $n_{\text{th}}^{*}$ increases in steps of one until it attains the maximum resolution of the PNR detector. This occurs because the Fock-basis population of the thermal-noise-perturbed state $\zeta_{0}^{\prime}$ still decreases monotonically with increasing photon number. This behavior differs from that under phase-diffusion noise (see Fig. 8(d)).
We then introduce the maximum tolerable thermal noise $n_{t}^{\text{max}}$ as a figure of merit, defined as the highest noise level for which $P_{\text{err}}^{\text{DSR}}\leq P_{\text{SQL}}^{\text{DSS}}$ at a given signal energy N. This quantity is shown in Fig. 10(e). Thus, the DSR outperforms the SQL when $n_{t}<n_{t}^{\text{max}}$, corresponding to the region below the $n_{t}^{\text{max}}$ curve. The plot displays M distinct peaks, after which it decays toward zero.
V Conclusion
We presented a displacement-squeeze receiver for S-BPSK discrimination based on Gaussian preprocessing and PNR/MAP thresholding. In the ideal, equal-prior case, $n_{\text{th}}^{*}=1$ and $P_{\text{err}}^{\text{DSR}}=\frac{1}{2}\exp\left[-4N\left(N+1\right)\right]$, independent of the PNR number resolution. The resulting performance is provably within a constant factor of the S-BPSK Helstrom bound for all N, $P_{\text{err}}^{\text{DSR}}\in\left[P_{\text{HB}}^{\text{DSS}},2P_{\text{HB}}^{\text{DSS}}\right]$, while surpassing the S-BPSK SQL at $N\approx 0.3$ and the C-BSPK Helstrom bound at $N\approx 0.4$ (sub-1% errors near $N\approx 0.6$). Under nonideal operation, we modeled detector inefficiency and dark counts, channel phase diffusion, and receiver thermal noise, and found that adaptive MAP thresholds maintain an advantage over the SQL within practical ranges (e.g., under $\sigma=0.1$, DSR exceeds the S-BPSK SQL for $N\in(0.28,0.93)$).
Future work includes extending the architecture to multi-ary constellations (e.g., QPSK and 16QAM) and validating performance in hardware under dynamic channel conditions.
Appendix A The Wigner Function of Displaced Squeezed Vacuum States
The Displaced Squeezed States (DSSs) discussed in this paper belong to single-mode Gaussian states, whose Wigner function is a bivariate Gaussian distribution fully characterized by two parameters: the displacement vector $\vec{d}$ and the covariance matrix $\mathbf{V}$. Given a DSS denoted as $\left|\psi\right\rangle$, its quantum mechanical expression is
$\left|\psi\right\rangle=\left|\alpha,r\right\rangle=D(\alpha)S(r)\left|0\right\rangle,\qquad\alpha,r\in\mathbb{R}_{+}.$ (40)
The first parameter, the displacement vector $\vec{d}$, is a two-dimensional vector representing the translation in phase space relative to the origin. Its components correspond to the expectation values of the quadrature operators $\hat{X},\hat{P}$:
$\vec{d}=\begin{pmatrix}d_{x}\\ d_{p}\end{pmatrix}=\begin{pmatrix}\left\langle\hat{X}\right\rangle\\[2.0pt] \left\langle\hat{P}\right\rangle\end{pmatrix}$. (41)
For $\left|\psi\right\rangle$, $d_{x}=\sqrt{2}\,\mathrm{Re}(\alpha)=\sqrt{2}\alpha$ and $d_{p}=\sqrt{2}\,\mathrm{Im}(\alpha)=0$. The second parameter, the covariance matrix $\mathbf{V}$, is a $2\times 2$ real symmetric positive-definite matrix describing the “shape” and “size” of the quantum state in phase space. Its elements are variances and covariances:
$\mathbf{V}$ $=\begin{bmatrix}V_{xx}&V_{xp}\\ V_{px}&V_{pp}\end{bmatrix}$ (42)
$=\begin{bmatrix}\left\langle(\Delta\hat{X})^{2}\right\rangle&\frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle\\[4.0pt] \frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle&\left\langle(\Delta\hat{P})^{2}\right\rangle\end{bmatrix}$,
where $\Delta\hat{O}=\hat{O}-\left\langle\hat{O}\right\rangle$, and the anticommutator is defined as $\left\{\hat{A},\hat{B}\right\}=\hat{A}\hat{B}+\hat{B}\hat{A}$. For $\left|\psi\right\rangle$, the covariance matrix is
$\mathbf{V}=\begin{bmatrix}\frac{1}{2}e^{-2r}&0\\[2.0pt] 0&\frac{1}{2}e^{2r}\end{bmatrix}$, (43)
indicating that in phase space, the state is an ellipse (coherent states and vacuum states appear as circles), with fluctuations squeezed along the $X$-direction and expanded along the $P$-direction, while the overall phase-space area remains unchanged ($\det\mathbf{V}=\tfrac{1}{2}e^{-2r}\cdot\tfrac{1}{2}e^{2r}=\tfrac{1}{4}$).
Substituting $\vec{d}$ and $\mathbf{V}$ into the general formula yields the Wigner function for $\left|\psi\right\rangle$:
$\begin{aligned} &W(\vec{r}|\psi)=\frac{1}{2\pi\sqrt{\det\mathbf{V}}}\,\exp\left\{-\frac{1}{2}(\vec{r}-\vec{d})^{T}\mathbf{V}^{-1}(\vec{r}-\vec{d})\right\}\\ &\to W(x,p|\psi)=\frac{1}{\pi}\exp\left\{-\,\mathrm{e}^{2r}\big(x-\sqrt{2}\alpha\big)^{2}-\,\mathrm{e}^{-2r}p^{2}\right\},\end{aligned}$ (44)
where $\vec{r}=\begin{pmatrix}x\\ p\end{pmatrix}$ is a point in phase space, and $\det\mathbf{V}$ denotes the determinant of $\mathbf{V}$.
Appendix B The Helstrom bound and SQL of S-BPSK under phase diffusion condition
Phase diffusion noise is a communication channel noise, therefore theoretically both the Helstrom bound and the SQL for S-BPSK $\rho_{i}^{\text{pd}}$ (Eq. (28)) are affected. Specifically, the SQL expression in Eq. (4) should be rewritten as:
$P_{\text{SQL}}^{\text{DSS}}\left(N,\sigma\right)$ $=\int_{0}^{+\infty}\mathrm{d}x\cdot P\left(x|\rho_{i}^{\text{pd}}\right)$ (45)
$=\int_{0}^{+\infty}\mathrm{d}x\int_{\mathbf{R}}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot P\left(x|\rho_{i,\phi}\right)$
$=\int_{0}^{+\infty}\mathrm{d}x\int_{\mathbf{R}}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\int_{\mathbf{R}}\mathrm{d}p\cdot W\left(x,p|\rho_{i,\phi}\right)$
$=\int_{\mathbf{R}}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot P_{\text{SQL}}^{\text{DSS}}(\phi)$
Here,
$P_{\text{SQL}}^{\text{DSS}}(\phi)=\int_{0}^{+\infty}\mathrm{d}x\cdot P\left(x|\rho_{i,\phi}\right)=\int_{0}^{+\infty}\mathrm{d}x\int_{\mathbf{R}}\mathrm{d}p\cdot W\left(x,p|\rho_{i,\phi}\right)$ (46)
represents the SQL for the binary quantum states $\left\{\rho_{0,\phi},\rho_{1,\phi}\right\}$ (see Eq. (29)). $P(x|\rho_{i,\phi})$ is the homodyne probability of obtaining outcome $x$ given the state $\rho_{i,\phi}$. Note: The rotation operator $R_{\phi}$ is a Gaussian operator, thus $\rho_{i,\phi}$ remains a Gaussian state. Its Wigner function can be computed using the method described in Appendix A.
For the Helstrom limit of two states $\rho_{i}^{\text{pd}}$, Eq. (2) (applicable only to pure states) is no longer suitable and should be replaced with its equivalent form:
$P_{\text{HB}}^{\text{DSS}}\left(N,\sigma\right)=\frac{1}{2}\left(1-\left\|P_{0}\rho_{0}^{\text{pd}}-P_{1}\rho_{1}^{\text{pd}}\right\|_{1}\right)$ (47)
where $\left\|A\right\|_{1}=\mathbf{Tr}\left\{A^{\dagger}A\right\}$ denotes the trace norm, and $P_{0},P_{1}$ denote prior probabilities. This expression applies to the calculation of the Helstrom bound for binary states, whether pure or mixed.
Acknowledgements.
This work was supported by Independent Innovation Science Fund Program of National University of Defense Technology, China (Grant No. 22-ZZCX-036), Key research&development program of Guangxi (Grant No. GuiKeAB23075155), the National Science Foundation for Young Scientists of China (Grant No. 12204539), and the Key Research & Development Program of Guangxi (Grant No. AB23075112).
Data Availability
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.